OpenStudy (anonymous):

C(x)=150+2250x-0.02x^2 find the marginal cost function

OpenStudy (rulnick):

C'(x) = 2250 - 0.04x

OpenStudy (anonymous):

can I know how you got that?

OpenStudy (rulnick):

take the derivative of the cost function to get the marg cost

OpenStudy (anonymous):

can I see the work?

OpenStudy (rulnick):

d(2250x)/dx = 2250 d(.02x^2)/dx = .04x that's about it